OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that models the values in the table? X | -2 | 0 | 4 F(x) | -7 | 3 | -73

OpenStudy (aum):

The standard form of a parabola is: y = ax^2 + bx + c when x = -2, y = -7. Put those values in y = ax^2 + bx + c and you will get one equation. Also put the other two points: (0,3) and (4, -73). You will have a total of three equations and three unknowns a, b and c. Solve the simultaneous equations for a,b and c.